# FAST: An Efficient Scheduler for All-to-All GPU Communication

### v2

@@ 4 Two Phase Scheduler
This section details how FAST mitigates GPU-level skew within each server through intra-server rebalancing (§4.1); addresses inter-server skew and incast by generating balanced, one-to-one transfer stages (§4.2); and integrates these mechanisms into a pipeline that overlaps intra- and inter-server traffic to optimize end-to-end performance (§4.3). We conclude by analyzing FAST’s key properties in §4.4.
4.1 Intra-server Scheduling: Balancing and Redistributing
The first challenge arises within each server: GPUs often generate or absorb uneven volumes of traffic, creating stragglers where some NICs sit idle while others are overloaded. By leveraging the ultra-fast scale-up fabric, our goal is to eliminate these sender- and receiver-side imbalances, presenting the scale-out tier with a more uniform workload (Figure˜6).
@@ 4.1 Intra-server Scheduling: Balancing and Redistributing
Mitigating sender skew. The first step is to prevent a GPU from being the ‘slow sender.’ In the tile, GPU must send 8 units, while only 4. If transmitted directly, would finish early, leaving as a straggler. To avoid this, we rebalance within server : heavily loaded GPUs shift part of their traffic to lightly loaded ones using the scale-up fabric. Here, transfers 2 units to , so both end up with 6. In matrix terms, the row sums of the tile are equalized—ensuring every NIC in contributes the same total outgoing load to server .
Mitigating receiver skew. After sender-side balancing, GPU may still receive 8 units (column sum) while only 4, leaving a receiver-side straggler. The fix is to decouple the notions of ‘correct server’ and ‘correct GPU.’ Each sender forwards all of its traffic to its peer GPU with the same local index (, ), ensuring data first arrives at the correct server. This merged peer transfer keeps receiver loads balanced—since senders were equalized earlier—even though some data arrives temporarily at the wrong GPU, to be corrected later. In matrix form, each row collapses into a single nonzero in its peer column, turning the tile into a scalar matrix with equal diagonal entries and all off-diagonals zero (right in Figure˜7). The result is one-to-one, balanced scale-out transfers across GPUs.
@@ 4.2 Inter-server Scheduling: Balanced, One-to-one Transfers
Although intra-server scheduling reshapes the initial skewed GPU-level All-to-All into a more balanced form, balancing over scale-up network within each server cannot eliminate server-to-server skew. Some servers still send or receive more traffic than others, creating bottlenecks that must be addressed to achieve good end-to-end performance.
This imbalance becomes clearer once we reduce the GPU-to-GPU traffic matrix into a simpler server-to-server view. Figure˜8 shows a 3-server, 2-GPU-per-server example: the original 6×6 GPU matrix (left) is reshaped by intra-server scheduling into the balanced form (middle), where each 22 server-to-server tile becomes a scalar matrix. Each scalar tile can then be collapsed into a single entry, yielding the reduced 33 server-level matrix (right). The intuition is that, after intra-server scheduling, GPUs within a server act identically over scale-out—each sending and receiving equal volumes—so we can abstract away individual GPUs. This reduction both exposes the remaining skew across servers and simplifies scheduling, since the server-level problem is typically an order of magnitude smaller than the GPU-level one (e.g., 8 GPUs per server in modern clusters).
@@
SpreadOut: one-to-one but not optimal. A natural incast-free baseline is MPI’s SpreadOut algorithm [33], which cycles through ‘shifted diagonals’ of the server matrix: at stage , server sends to server . This guarantees one-to-one sender-receiver mappings at every stage.
However, SpreadOut fails the second requirement: the bottleneck server may sit idle during many stages. Although the bottleneck is the row or column with the largest sum, its per-stage assignment may not be the largest entry on that diagonal. When this happens, the stage is gated by another flow elsewhere, forcing the true bottleneck to wait.
Figure˜9 shows an example: server , as a receiver is the bottleneck—with 14 units column sum—the heaviest among all rows/columns. But in stage 1, receives only 3 units, while the stage is gated by a separate 5-unit flow ( where acts as a sender). As a result, sits idle as a receiver for 2 extra time units. The same situation recurs in stage 3, adding another 1 unit of idle time. Altogether, SpreadOut finishes in 17 units—3 units slower than the 14-unit theoretical minimum.
@@
Birkhoff’s decomposition [8] is tailored to this setting. It expresses any traffic matrix666Formally, the theorem applies to scaled doubly stochastic matrices; arbitrary matrices can be adapted as described later. as a weighted sum of permutation matrices. Each permutation matrix has exactly one nonzero per row and column, all of identical value, representing a one-to-one, balanced transfer stage: every active sender transmits the same amount to exactly one receiver, and each receiver accepts from one sender. Some permutation matrices may be partial, with entire rows or columns set to zero, corresponding to servers that have already finished. In Figure˜9, the first four stages are full permutation matrices, while the last two are partial.
@@
Viewed as a schedule, this decomposition yields a sequence of one-to-one server-receiver matchings where bottleneck servers remain continuously active until they complete. In our example, Birkhoff finishes in 14 units—exactly the lower bound—achieving optimality.
@@
Multi-server end-to-end scheduling. The final piece of our scheduler—Birkhoff’s decomposition—addresses the remaining server-level skew. We illustrate the full scheduling process with a 3-server, 2-GPU-per-server example in Figure˜10, omitting intra-server transfers (grey diagonal tiles) for clarity. The input is a 66 GPU-to-GPU traffic matrix (left). If GPUs transmit directly without our scheduler, the completion time lower bound is 10 time units, dictated by the heaviest sender GPU (, row sum 10) and busiest receiver GPU (, column sum 10).
@@
Step 1: Balancing. This step reduces the severity of the bottleneck. Within each 22 tile (e.g., ), sender loads are equalized across GPUs, and peer transfer () ensures receivers share the load evenly. With intra-server skew removed, the effective lower bound improves: in the reshaped matrix (middle of Figure˜10), the maximum row/column sum drops from 10 to 8 ( as sender and as receiver). Intuitively, the pressure of a straggling NIC/GPU is averaged across all GPUs and NICs within a server, reducing its impact. At this point, the 66 GPU-level matrix can be cleanly collapsed into a skewed 33 server-level alltoallv.
@@
Step 2: Balanced, one-to-one transfer stages. Birkhoff’s decomposition then partitions this server-level matrix into three balanced, one-to-one transfer stages (right). Each stage delivers a portion of the workload, and together they complete all transfers. The resulting schedule satisfies three key properties: (i) Incast-free: At the server level, Birkhoff enforces one-to-one matchings. Combined with Step 1’s peer-access rule, each GPU communicates only with the same-index GPU in the matched server, preventing any receiver overload. (ii) Balanced: Servers send equal volumes per stage, while Step 1 guarantees balanced GPUs within each server. (iii) Optimal: Bottleneck servers (e.g., as senders as receiver here) remain fully active across all stages, achieving the theoretical minimum completion time (8 units).
Step 3: Per-stage redistribution. Step 2’s scale-out transfers ensure data reaches the correct server, but not necessarily the correct GPU. A lightweight redistribution step fixes placement locally, aligned with each stage. For example, in Figure˜10, once (blue tile) completes in Stage1, the corresponding portion is immediately redistributed within (shown as the blue-striped tile).
@@ 4.3 End-to-End Transfer Pipeline
So far, our scheduler combines scale-up transfers (balancing and redistribution) with scale-out transfers (staged by Birkhoff’s decomposition). While scale-up is much faster than scale-out, simply serializing everything, e.g., balance stage 1 scale-out stage 1 redistribute stage 2 scale-out , still results in noticeable overhead. To minimize end-to-end completion time, we build a pipeline that keeps the scale-out network—the true bottleneck—as busy as possible, while pushing scale-up operations into the background.
There are three types of scale-up transfers to consider: (i) Balancing, which equalizes GPU loads before scale-out begins; (ii) Redistribution, which was a one-time step in intra-server scheduling (§4.1) but is now aligned per stage, correcting placement as each stage’s scale-out completes (striped submatrices in Figure˜10); and (iii) Intra-server portion of alltoallv (grey diagonal tiles in Figure˜10). The scale-out portion is straightforward: successive one-to-one stages from Birkhoff’s decomposition.
Pipeline structure. Figure˜11 illustrates how the pipeline operates. Balancing (purple) runs first, as all subsequent scale-out stages depend on the reshaped workload. Once balancing finishes, scale-out begins immediately. During scale-out, Stage ’s redistribution (blue) overlaps with Stage ’s scale-out (green), hiding almost all redistribution cost—except in the final stage, which has no successor. The intra-server portion of alltoallv (grey) is executed alongside the first scale-out stage, making use of idle scale-up bandwidth before redistributions are triggered. While the pipeline could be made even tighter by subdividing balancing and scale-out into smaller chunks, the gain is marginal, so we adopt this simpler design.
Overall, the pipeline keeps the scale-out network—the true bottleneck—fully utilized, while scale-up operations are largely hidden in the background. This completes the design: intra-server scheduling removes GPU-level skew, inter-server scheduling produces balanced server-level transfers, and pipelining integrates them into a seamless end-to-end execution. Next, we turn to the key properties of our scheduler, including optimality and computational complexity.
@@ 4.4 Scheduler Properties
Beyond the algorithm itself, the scheduler’s properties reveal why it is a practical, efficient, and high-performance solution for alltoallv in modern GPU clusters.
@@
Optimality. The majority of end-to-end transfer—staged scale-out transfers generated by Birkhoff’s decomposition—operates at full efficiency on the scale-out fabric. Suboptimality arises from additional intra-server operations such as balancing and redistribution. Empirically, these costs are small: under typical workloads, they add less than 5% to the total scale-out cost (as shown in Section˜5.1.3), and even under highly skewed traffic (Zipfian with skewness 0.95), the overhead remains below 8%.
@@
We also prove in Appendix A.1 that even under adversarial workloads—designed to maximize balancing (e.g., all GPUs within a server join All-to-All but only one holds data) and redistribution (e.g., only one GPU in each destination receives data)—the performance gap from the theoretical optimum remains bounded. For instance, in a 4-node cluster with a 9:1 scale-up to scale-out bandwidth ratio (450 GBps NVLink [37] vs. 400 Gbps Ethernet), FAST finishes within 2.12 of the optimum, even in this worst-case setting.
@@
Number of stages. The number of transfer stages produced by Birkhoff’s decomposition depends on the server-level matrix. In the best case, a perfectly balanced matrix requires exactly stages—the same as SpreadOut. With skew, more stages may be introduced, but the total is always bounded by [22].
@@
Fewer stages are preferable because each adds synchronization overhead. Ideally, one would minimize the stage count, but finding such a decomposition is NP-hard [15]. Rather than attempting this costly search, our scheduler efficiently produces a valid decomposition. Since the stage count is bounded, the synchronization cost is also bounded—and in practice, our evaluation shows it to be negligible.
@@
Computational complexity. Intra-server scheduling is lightweight, involving only simple load balancing and redistribution bookkeeping. The main cost lies in inter-server scheduling—Birkhoff’s decomposition—which runs in polynomial time, with complexity. Here, is the number of servers, not GPUs, since decomposition is applied only to the reduced server-level matrix.
@@
In practice, scheduling is extremely fast on modern CPUs.777Figure 12 measured on Intel Xeon Platinum 8468. As shown in Figure˜12, our scheduler completes in 25 µs for 4 servers (32 GPUs), 221 µs for 8 servers (64 GPUs), and 805 µs for 12 servers (96 GPUs), assuming 8 GPUs per server—a common configuration. With one expert per GPU, this covers the typical expert parallelism (EP) range of today’s MoE workloads: from EP8–EP128 (e.g., Perplexity AI [6]) up to very large deployments like EP320 (DeepSeek [14]). Even at EP320—40 servers—the runtime is only 77 ms, well within practical bounds.
@@
To put this in context, consider a median-scale case with EP64: each GPU transmits about 1 GB of data to others—the data exchange scale reported by prior work [25, 46]. Over a 400 Gbps network, such an All-to-All takes at least 20 ms, while scheduling adds just 221 s (1.1% of total time). Our scheduling step is a small upfront ‘tax’ that yields a fully optimized plan, shortening end-to-end completion compared to unscheduled transfers.
@@
By contrast, prior approaches [44, 27, 10] cast scheduling as NP-hard problems (e.g., MILP or multi-commodity flow) and rely on solvers [18] that run orders of magnitude slower: for a 16-GPU All-to-All, the fastest solver-based scheduler, SyCCL, takes 3.6 s—while our scheduler runs in 3.1 s.
Adaptation of an arbitrary input matrix to a valid form. In theory, Birkhoff’s theorem applies to scaled doubly stochastic matrices, where every row and column sum is equal. Since real server matrices are arbitrary, we first embed them into this form by constructing an auxiliary matrix (always possible in ). The procedure increases only the lighter rows or columns until all sums match the heaviest one, leaving the bottleneck row/column unchanged.
This step is purely mathematical: the auxiliary entries represent ‘virtual’ transfers that are never executed. In practice, once all real traffic has finished, the auxiliary load is ignored. This is why some permutation matrices produced by the decomposition appear partial (Figure˜9): rows or columns of all zeros stem from the auxiliary matrix. Importantly, this conversion preserves both correctness and optimality, as the true bottleneck—the largest row or column sum—remains unchanged.
@@
Exclusion of All-to-All scheduling over scale-up. Both balancing and redistribution are themselves skewed alltoallv operations. Because they run entirely on the high-bandwidth scale-up fabric, sophisticated scheduling is unnecessary. For these steps we use MPI’s SpreadOut algorithm [33], which provides simple one-to-one sender–receiver mappings at low cost. While Birkhoff’s decomposition could deliver an optimal schedule here as well, its added computation is unnecessary—the scale-up tier is not the bottleneck, so the performance difference is negligible.
@@
Another caveat is that SpreadOut may not be well suited for older GPUs with non-symmetric scale-up topologies, such as the ring in AMD MI250 [1] or the hybrid cube mesh in NVIDIA V100 [36]. However, recent GPUs adopt mostly symmetric intra-server configurations, e.g., the switch-based topology [37] and the fully connected mesh [2], for scale-up fabric, which are our primary target platforms.

### v3

@@ 4 Two Phase Scheduler
This section presents our design: mitigating skew within servers (§4.1), handling incast and imbalance across servers (§4.2), and pipelining both to improve end-to-end transfers (§4.3). We conclude with an analysis of key properties (§4.4).
4.1 Intra-server Scheduling: Balancing and Redistribution
The first challenge arises within each server: GPUs often generate or absorb uneven volumes of traffic, creating stragglers where some NICs sit idle while others are overloaded. By leveraging the fast scale-up fabric, our goal is to eliminate these sender- and receiver-side imbalances, presenting the scale-out tier with a more uniform workload.
@@ 4.1 Intra-server Scheduling: Balancing and Redistribution
Mitigating sender skew. The first step is to prevent a GPU from being the ‘straggler sender’. In the tile, GPU must send 8 units, while only needs 4. If transmitted directly, would finish early, leaving as a straggler. To avoid this, we rebalance within server : heavily loaded GPUs shift part of their traffic to lightly loaded ones using the scale-up fabric. Here, transfers 2 units to , so both end up with 6. In matrix terms, the row sums of the tile are equalized—ensuring every NIC in contributes the same total outgoing load to server .
Mitigating receiver skew. After sender-side balancing, GPU may still receive 8 units (column sum) while only needs 4, leaving a receiver-side straggler. The fix is to decouple the notions of ‘correct server’ and ‘correct GPU’. Each sender forwards all of its traffic to its peer GPU with the same local index (, ), ensuring data first arrives at the correct server. This merged peer transfer keeps receiver loads balanced—since senders were equalized earlier—even though some data arrives temporarily at the wrong GPU, to be corrected later. In matrix form, each row collapses into a single nonzero in its row-local-index column, turning the tile into a scalar matrix with equal diagonal entries and all off-diagonals zero (right in Figure˜7). The result is one-to-one, balanced scale-out transfers across GPUs.
@@ 4.2 Inter-server Scheduling: Balanced, One-to-one Transfers
Although intra-server scheduling reshapes the initial skewed GPU-level alltoallv into a more balanced form, balancing over scale-up network within each server cannot eliminate server-to-server skew. Some servers still send or receive more traffic than others, creating bottlenecks that must be addressed to achieve good end-to-end performance.
This imbalance becomes clearer once we reduce the GPU-to-GPU traffic matrix into a simpler server-to-server view. Figure˜8 shows a 3-server, 2-GPU-per-server example: the original 6×6 GPU matrix (left) is reshaped by intra-server scheduling into the balanced form (middle), where each 22 server-to-server tile becomes a scalar matrix. Each scalar tile can then be collapsed into a single entry, yielding the reduced 33 server-level matrix (right). The intuition is that, after intra-server scheduling, GPUs within a server act identically over scale-out—each sending and receiving equal volumes—so we can abstract away individual GPUs. This reduction both exposes the remaining skew across servers and simplifies scheduling, since the server-level problem is typically an order of magnitude smaller than the GPU-level one (e.g., with 8 GPUs per server [hgx] in modern clusters).
@@
SpreadOut: one-to-one but not optimal. A natural incast-free baseline is MPI’s SpreadOut algorithm [mpi_all2all], which cycles through ‘shifted diagonals’ of the server matrix: at stage , server sends to server . This guarantees one-to-one sender-receiver mappings at every stage.
However, SpreadOut fails the second requirement: the bottleneck server may sit idle during many stages. Although the bottleneck is the row or column with the largest sum, the entry selected for that row or column in a given stage need not be the largest entry on the corresponding diagonal. When this happens, the stage is gated by another matrix entry (i.e., a flow from non-bottlenecks), forcing the true bottleneck to wait.
Figure˜9 shows an example: server , as a receiver, is the bottleneck—with 14 units column sum—the heaviest among all rows/columns. But in stage 1, receives only 3 units, while the stage is gated by a separate 5-unit flow ( where is a sender). As a result, sits idle as a receiver for 2 extra time units. The same situation recurs in stage 3, adding another 1 unit of idle time. Altogether, SpreadOut finishes in 17 units—3 units slower than the 14-unit theoretical minimum.
@@
Birkhoff’s decomposition [birkhoff] is tailored to this setting. It expresses a traffic matrix as a weighted sum of permutation matrices. Each permutation matrix has exactly one nonzero per row and column, all of identical value, representing a one-to-one, balanced transfer stage—every active sender transmits the same amount to exactly one receiver, and each receiver accepts from one sender. Some permutation matrices may be partial, with zero rows or columns for servers that have already finished. In Figure˜9, the first four stages are full permutation matrices, while the last two are partial.
@@
Viewed as a schedule, this decomposition yields a sequence of one-to-one sender-receiver matchings where bottleneck servers remain continuously active until they complete. In our example, Birkhoff finishes in 14 units—exactly the lower bound—achieving optimality.
@@
Multi-server end-to-end scheduling. The final piece of our scheduler—Birkhoff’s decomposition—addresses the remaining server-level skew. We illustrate the full scheduling process with a 3-server, 2-GPU-per-server example in Figure˜10, omitting intra-server transfers (grey diagonal tiles) for clarity. The input is a 66 GPU-to-GPU traffic matrix (left). Without our scheduler, the completion-time lower bound is 10 units, set by the heaviest sender GPU (, row sum 10) and receiver GPU (, column sum 10).
@@
Step 1: Balancing. This step reduces the severity of the bottleneck. Within each 22 tile (e.g., ), sender loads are equalized across GPUs, and peer transfer (e.g., ) ensures receivers share the load evenly. With intra-server skew removed, the effective lower bound improves: in the reshaped matrix (middle of Figure˜10), the maximum row/column sum drops from 10 to 8 (i.e., as sender and as receiver). Intuitively, the pressure of a straggling NIC/GPU is averaged across all GPUs and NICs within that server, reducing its impact. At this point, the 66 GPU-level matrix can be cleanly collapsed into a skewed 33 server-level alltoallv.
@@
Step 2: Balanced, one-to-one transfer stages. Birkhoff’s decomposition then partitions this server-level matrix into three balanced, one-to-one transfer stages (right). Each stage delivers a portion of the workload, and together they complete all transfers. The resulting schedule satisfies three key properties: (i) Incast-free: At the server level, Birkhoff enforces one-to-one matchings. Combined with Step 1’s peer-access rule, each GPU communicates only with the same-index GPU in the matched server, preventing any receiver overload. (ii) Balanced: Servers send equal volumes per stage, while Step 1 guarantees balanced GPUs within each server. (iii) Optimal: Bottleneck servers (i.e., as senders, as receiver here) remain fully active across all stages, achieving the theoretical minimum completion time (8 units).
[…]
@@ 4.3 End-to-End Transfer Pipeline
Our scheduler combines scale-up transfers (balancing and redistribution) with scale-out transfers (staged by Birkhoff’s decomposition). While scale-up is much faster than scale-out, serializing all the steps, e.g., balance stage 1 scale-out stage 1 redistribute stage 2 scale-out , still results in noticeable overhead. To minimize end-to-end completion time, we build a pipeline that keeps the scale-out tier—the true bottleneck—as busy as possible, while hiding scale-up transfers in the background.
There are three types of scale-up transfers to consider: (i) Balancing, which equalizes GPU loads before scale-out begins; (ii) Redistribution, which corrects placement as scale-out completes, now aligned per stage (e.g., the striped tiles in Figure˜10); and (iii) Intra-server portion of alltoallv (grey tiles in Figure˜10). The scale-out portion is straightforward: successive one-to-one stages from Birkhoff’s decomposition.
Pipeline structure. Figure˜11 illustrates how the pipeline operates. Balancing (purple) runs first, as all subsequent scale-out stages depend on the reshaped workload. Once balancing finishes, scale-out begins immediately. During scale-out, Stage ’s redistribution (blue) overlaps with Stage ’s scale-out (green), hiding redistribution cost—except in the final stage, which has no successor. The intra-server portion of alltoallv (grey) is executed alongside the first scale-out stage, making use of idle scale-up bandwidth before redistributions are triggered. While the pipeline could be made even tighter by subdividing balancing and scale-out into smaller chunks, the gain is small, so we adopt this simple design.
Overall, the pipeline keeps the scale-out network—the true bottleneck—as busy as possible, while scale-up operations are largely hidden in the background. This completes the design: intra-server scheduling removes GPU-level skew, inter-server scheduling produces balanced server-level transfers, and pipelining integrates them into a seamless end-to-end execution. Next, we turn to the key properties of our scheduler, including optimality and computational complexity.
@@ 4.4 Scheduler Properties
Beyond the algorithm itself, the scheduler’s properties reveal why it is a practical, efficient, and high-performance solution for alltoallv in modern GPU clusters. Assume servers and each server has GPUs.
@@
Optimality. The majority of end-to-end transfer—staged scale-out transfers generated by Birkhoff’s decomposition—operates at full efficiency on the scale-out fabric. Suboptimality arises from additional intra-server operations such as balancing and redistribution. Empirically, these costs are small: under typical workloads, they add less than 5% to the total scale-out cost (as shown in Section˜5.1.3), and even under highly skewed traffic (Zipfian with skewness 0.9), the overhead remains below 8%.
@@
We also prove in Appendix A.1 that even under adversarial workloads—designed to maximize balancing (e.g., all GPUs within a server join alltoallv but only one holds data) and redistribution (e.g., only one GPU in each destination server receives data)—the performance gap from the theoretical optimum remains bounded. For instance, in a 4-node cluster with a 9:1 scale-up to scale-out bandwidth ratio (450 GBps 4th-gen NVLink [h100] vs. 400 Gbps Ethernet), FAST finishes within 2.12 of the optimum, even in this worst-case setting.
@@
Number of stages. The number of transfer stages produced by Birkhoff’s decomposition depends on the server-level matrix. In the best case, a perfectly balanced server-level matrix requires exactly stages—the same as SpreadOut. With skew, more stages may be introduced, but the total is always bounded by [birkhoff_stages].
@@
Fewer stages are preferable because each adds synchronization overhead. Ideally, one would minimize the stage count, but finding such a decomposition is NP-hard [birkhoff_nphard]. Rather than attempting this costly search, our scheduler efficiently produces a valid decomposition. Since the stage count is bounded, the synchronization cost is also bounded—and in practice, our evaluation shows it to be negligible.
@@
Computational complexity & runtime. Intra-server scheduling is lightweight, involving only simple load balancing and bookkeeping. The primary cost lies in inter-server scheduling—Birkhoff’s decomposition—which runs in polynomial time with complexity and is fast in practice.
@@
As shown in Figure˜16, our scheduler completes in 25 µs for 4 servers (32 GPUs), 221 µs for 8 servers (64 GPUs), and 805 µs for 12 servers (96 GPUs), assuming GPUs per server—a common configuration [hgx]. With one expert per GPU, this spans the typical expert-parallelism (EP) range of today’s MoE workloads, from EP8—EP128 (e.g., Perplexity AI [perplexityai]) to large-scale deployments such as EP320 (DeepSeek [deepseek]). Even at EP320 (40 servers), scheduling overhead remains modest at 77 ms, well within practical bounds.
@@
To put this in context, consider a median-scale case with EP64: each GPU transmits about 1 GB of data to others—the traffic scale reported by prior work [mixnet, deepep]. Over a 400 Gbps network, such an All-to-All takes at least 20 ms, while scheduling adds 221 s (1.1% of total time). Our scheduling step is a small upfront ‘tax’ that yields a fully optimized plan, shortening end-to-end completion compared to no schedules.
@@
By contrast, prior approaches [taccl, teccl, syccl] cast scheduling as NP-hard problems (e.g., MILP or multi-commodity flow) and rely on solvers [gurobi] that run orders of magnitude slower: for a 16-GPU All-to-All, the fastest solver-based scheduler, SyCCL [syccl], takes 3.6 s—while our scheduler runs in 3.1 s.
@@
Exclusion of All-to-All scheduling over scale-up. Both balancing and redistribution are themselves skewed alltoallv operations. Because they run entirely on the fast scale-up fabric, sophisticated scheduling is unnecessary. For these steps we use MPI’s SpreadOut algorithm [mpi_all2all], which provides simple one-to-one sender–receiver mappings at low cost. While Birkhoff’s decomposition could deliver an optimal schedule here as well, its added computation is unnecessary—the scale-up tier is not the bottleneck.
@@
Another caveat is that SpreadOut may not be well suited for older GPUs with non-symmetric scale-up topologies, e.g., the ring in AMD MI250 [mi250x] and the hybrid cube mesh in NVIDIA V100 [v100]. However, recent GPUs adopt symmetric scale-up fabric, e.g., switch-based topology [h100] and fully connected mesh [mi300x], which are our target platforms.
Adapting an arbitrary matrix to a valid form. Birkhoff’s theorem applies to scaled doubly stochastic matrices, where all row and column sums are equal. Since real server-level traffic matrices are arbitrary, we first embed them into this form by adding an auxiliary matrix, which can be constructed in time. This procedure increases only the lighter rows or columns until all sums match the heaviest one, leaving the true bottleneck row or column unchanged.
The auxiliary entries represent virtual transfers that are never executed and are ignored once all real traffic completes. As a result, some permutation matrices produced by the decomposition may appear partial with respect to real traffic (Figure˜9), with zero rows or columns originating from the auxiliary matrix. Importantly, this transformation preserves both correctness and optimality, since the maximum row or column sum—the true bottleneck—remains unchanged.
Birkhoff’s decomposition: overview. The algorithm takes as input an scaled doubly stochastic matrix and outputs a sequence of permutation matrices whose weighted sum reconstructs the input. The matrix can be viewed as a bipartite graph with senders (rows) and receivers (columns), where nonzero entries correspond to edges.
At a high level, the algorithm repeatedly finds a perfect matching in this graph; each such matching selects exactly one outgoing edge per sender and one incoming edge per receiver, yielding a permutation matrix. Each matching can be computed, for example, using the Hungarian algorithm [Kuhn1955Hungarian] with complexity. After subtracting the permutation matrix derived from each matching, the residual matrix remains scaled doubly stochastic, allowing the process to repeat. In the worst case, the algorithm requires iterations [birkhoff_stages], resulting in total complexity.
A key advantage of the algorithm is that it advances all bottleneck rows and columns—potentially multiple with equal maximum load—at the same rate. In contract, a greedy algorithm may fail to account for all bottlenecks simultaneously, often prioritizing individual large entries and suboptimal.